# Peri-LN: Revisiting Layer Normalization in the Transformer Architecture

### v2

@@ Abstract
Designing Transformer architectures with the optimal layer normalization (LN) strategy that ensures large-scale training stability and expedite convergence has remained elusive, even in this era of large language models (LLMs). To this end, we present a comprehensive analytical foundation for understanding how different LN strategies influence training dynamics in large-scale Transformer training. Until recently, Pre-LN and Post-LN have long dominated standard practices despite their limitations in large-scale training. However, several open-source large-scale models have recently begun silently adopting a third strategy without much explanation. This strategy places layer normalization (LN) peripherally around sublayers, a design we term Peri-LN. While Peri-LN has demonstrated promising empirical performance, its precise mechanisms and benefits remain almost unexplored. Our in-depth analysis shows that Peri-LN strikes an ideal balance in variance growth—unlike Pre-LN and Post-LN, which are prone to vanishing gradients and “massive activations.” To validate our theoretical insight, we conduct large-scale experiments on Transformers up to $3.2$B parameters, showing that Peri-LN consistently achieves more balanced variance growth, steadier gradient flow, and convergence stability. Our results suggest that Peri-LN warrants broader consideration for large-scale Transformer architectures, providing renewed insights into the optimal placement and application of LN.
@@ 1 Introduction
Building on a rapidly expanding lineage of Transformer-based large language models, open-source models have shown remarkable impact (Hoffmann et al., 2022; Dubey et al., 2024; Guo et al., 2025). As the demand for larger and more powerful models grows, various training stabilization techniques have been introduced (Yang et al., 2022; Zhai et al., 2023; Loshchilov et al., 2024). Among these, the choice of where and how to apply layer normalization (LN: LayerNorm or RMSNorm; Ba et al., 2016; Zhang & Sennrich, 2019) critically influences model convergence (Xiong et al., 2020; Kedia et al., 2024; Wortsman et al., 2024; Chung et al., 2024). However, their immense computational requirements have restricted deeper exploration of the underlying Transformer structure. Are we truly employing the optimal LN placement? In practice, fully revealing the results of massive resource investments can be challenging (Rivière et al., 2024). Despite its importance, there is still no consensus on a single best LN placement strategy.
@@
Two prominent LN placements have been widely explored. Post-LN (Vaswani et al., 2017) normalizes the hidden state after adding the sub-layer output to the residual stream (that is, $\mathrm{Norm}(x+\mathrm{Module}(x))$ where $x$ is input hidden state. $\mathrm{Norm}$ is LN). This helps constrain the variance of hidden states but may inadvertently weaken gradient signals, particularly in deeper models (Kedia et al., 2024). Pre-LN (Dubey et al., 2024), by contrast, normalizes before passing the hidden state to the sub-layer (that is, $x+\mathrm{Module}(\mathrm{Norm}(x))$). While this can enhance gradient propagation, it also admits so-called “massive activations,” where hidden states grow exponentially across layers (Sun et al., 2024; Wortsman et al., 2024; Zhai et al., 2023).
@@
Previous studies on deep convolutional neural networks (CNNs) have analyzed the impact of batch normalization on variance changes during the initialization stage of ResNet architectures, demonstrating its relationship to model performance (De & Smith, 2020). They noted that, in models without normalization, hidden activation growth at initialization can be exponential, leading to poor performance and stability. In contrast, in pre-normalized CNNs, the variance of hidden activations was shown to increase linearly as model depth grows. In the same vein, Kedia et al. (2024) reported that, for Transformer architectures as well, the variance in the forward propagation of Transformer-based language models at initialization increases linearly with depth. However, in the context of Transformer architectures, we observed that this variance growth at initialization does not persist as training progresses as shown in Figure 1. Section 3 and 4 provide a more detailed discussion of these hidden-state growth patterns.
@@
Beyond these two common strategies, Post-LN and Pre-LN, a third LN placement has quietly emerged in large-scale open-source models: applying LN around the sub-layer, i.e., on both its input and output. Although recent open-source models (Rivière et al., 2024; OLMo et al., 2024) have quietly adopted such designs and demonstrated promising performance on a large scale, these efforts often appeared isolated, lacking a conceptual unifying framework or a thorough investigation into their benefits. In this paper, we coin the term Peri-LN111“Peri-” means “around,” reflecting that LN encapsulates the entire sub-layer. to unify these scattered approaches and highlight an underexplored avenue for stabilizing large-scale Transformer training. While prior work offers encouraging empirical results, there is still a lack of in-depth analysis explaining why and how Peri-LN can help mitigate issues that frequently arise with Post-LN and Pre-LN. By probing the forward and backward dynamics of Peri-LN, we provide fresh insights into where and when it may offer advantages over more widely adopted LN placements.
@@
    Present an in-depth empirical analysis of Post-LN and Pre-LN in large-scale Transformers, examining how variance and gradient properties evolve beyond initialization. By studying forward and backward propagation statistics, we clarify why Post-LN may suffer vanishing gradients and why Pre-LN can exhibit exponential growth of hidden state.
@@
    Investigate Peri-LN to understand how normalizing both the inputs and outputs of each module moderates hidden-state behavior during forward and backward propagation, providing a systematic perspective on this underexplored alternative. We show that Peri-LN not only improves gradient stability and final loss but also plays a critical role in reducing hidden-state redundancy, as demonstrated by our ablation studies.
@@
    Provide quantitative evidence on how large activation influences training stability, benchmark performance, and redundancy in hidden states. Drawing on theoretical bounds for gradient propagation as well as extensive experiments, we demonstrate that Peri-LN can mitigate excessive variance growth and gradient instability more effectively than commonly used alternatives.
Figure 1: Illustration of hidden-state variance across different model depths and training iterations. From the initialization stage up to the point where 6.3 billion tokens were trained, we observed the variance growth of hidden states for Pre-LN and Post-LN architectures. The analysis was conducted using a $1.5$B-parameter model, and consistent trends were observed across models of different sizes. Detailed settings and more results are in Section 4.4.2.
@@ 2 Background and Motivation
The analysis of activation variance at model initialization has long been central to understanding normalization layers and enhancing stability in convolutional neural networks (CNNs) (De & Smith, 2020; He et al., 2016; Brock et al., 2021a). De & Smith (2020) showed that batch normalization in residual blocks can bias networks toward the identity function, thereby stabilizing gradients and improving overall training dynamics.
Similar investigations have emerged for Transformer architectures, examining how variance propagates and how gradients behave in both post-layer normalization (Post-LN) (Vaswani et al., 2017) and pre-layer normalization (Pre-LN) (Dubey et al., 2024) configurations (Xiong et al., 2020; Kedia et al., 2024; Wortsman et al., 2024; Li et al., 2024b). Early work comparing Post- and Pre-LN primarily focused on gradient scales and loss behavior. Xiong et al. (2020) observed that Pre-LN architectures tend to exhibit more stable gradients, but can still encounter issues such as gradient spikes and divergence, especially in deeper models or large-scale pre-training scenarios (Zhai et al., 2023; Wortsman et al., 2024; Fishman et al., 2024; Chung et al., 2024).
Among these challenges, the phenomenon of “massive activations” has attracted particular attention (Dettmers et al., 2022; Yu et al., 2024; Fishman et al., 2024). Sun et al. (2024) identified that in Pre-LN architectures, large spikes in activation magnitude can persist across layers due to residual connections. These massive activations act as fixed biases, potentially narrowing the model’s focus to certain tokens and may influence generalization. However, the underlying mechanisms behind these large values—and their exact impact on the training process—remain not yet well understood.
@@
These observations underscore the ongoing need to clarify how activation dynamics, normalization strategies, and architectural choices interact, especially in large-scale models. In response, this work aims to deepen our understanding of activation evolution during Transformer training under different normalization architectures, focusing on the role of massive activations and their effects on overall stability and performance.
We defer an extended discussion of the related literature to Appendix A, owing to space limitations.
3 Comparative Analysis
In this section, we discuss how different placements of layer normalization (LN  [^2] Unless stated otherwise, LN refers to both LayerNorm and RMSNorm.) in Transformer architecture affect both training stability and the statistics of hidden states (activations  [^3] We use “hidden state” and “activation” interchangeably.).
@@ A B C
Table 1: Intuitive comparison of normalization strategies.
Strategy Variance Growth Gradient Stability
Post-LN Mostly constant Potential for vanishing
Pre-LN Exponential in depth Potential for explosion
[…]
@@ 3.2 Variance Behavior from Initialization to Training
We find that Post-LN maintains a roughly constant variance, which helps avert exploding activations. Yet as models grow deeper and training proceeds, consistently normalizing $x_{l}+\mathrm{Module}(x_{l})$ can weaken gradient flow, occasionally causing partial vanishing gradients and slower convergence. In contrast, Pre-LN normalizes $x_{l}$ before the module but leaves the module output unnormalized, allowing hidden-state variance to accumulate exponentially once parameter updates amplify the input. Although Pre-LN preserves gradients more effectively in earlier stages, this exponential growth in variance can lead to “massive activations” (Sun et al., 2024), risking numeric overflow and destabilizing large-scale training. We reconfirm this in Section 4.
Takeaways.
  • Keeping the Highway Clean: Post-LN’s Potential for Gradient Vanishing and Slow Convergence. When layer normalization is placed directly on the main path (Placement $C$ in Figure 2), it can cause gradient vanishing and introduce fluctuations in the gradient scale, potentially leading to instability.
  • Maintaining a Stable Highway: Pre-LN May Not Suffice for Training Stability. Pre-LN does not normalize the main path of the hidden states, thereby avoiding the issues that Post-LN encounters. Nevertheless, a structural characteristic of Pre-LN is that any large values arising in the attention or MLP modules persist through the residual identity path. In particular, as shown in Figure 1, the exponentially growing magnitude and variance of the hidden states in the forward path may lead to numerical instability and imbalance during training.
Recent open-sourced Transformer architectures have adopted normalization layers in unconventional placements. Models like Gemma2222 and OLMo2222 utilize normalization layers at the module output (Output-LN), but the benefits of these techniques remain unclear (Rivière et al., 2024; OLMo et al., 2024). To investigate the impact of adding an Output-LN, we explore the peri-layer normalization architecture.
3.3 Placing Module Output Normalization
@@
    Initial Embedding Normalization:
@@
    $y_{l}=x_{l}+\mathrm{Norm}\Bigl{(}\mathrm{Module}\bigl{(}\mathrm{Norm}(x_{l})\bigr{)}\Bigr{)},$
@@
By normalizing both the input and output of each sub-layer, Peri-LN constrains the residual spikes common in Pre-LN, while retaining a stronger gradient pathway than Post-LN. Concretely, if $\mathrm{Norm}(\mathrm{Module}(\mathrm{Norm}(x_{l})))$ has near-constant variance $\beta_{0}$, then
$\mathrm{Var}(x_{l+1})\;\approx\;\mathrm{Var}(x_{l})+\beta_{0},$
leading to linear or sub-exponential hidden state growth rather than exponential blow-up. We empirically verify this effect in Section 4.4.2.
@@
Figure 3: Performance comparison of Post-LN, Pre-LN, and Peri-LN Transformers during pre-training. Figure 3(a) llustrates the pre-training loss across learning rates. Pre-training loss and gradient norm of best performing $400$M size Transformers are in Figure 3(b) and 3(c). Consistent trends were observed across models of different sizes.
(a) Training loss
(b) Loss in the final $5$B token interval
(c) Gradient-norm
Open-Sourced Peri-LN Models: Gemma2222 & OLMo2222.
Both Gemma2222 and OLMo2222, which apply output layer normalization, employ the same peri-normalization strategy within each Transformer layer. However, neither model rigorously examines how this placement constrains variance or mitigates large residual activations. Our work extends Gemma2222 and OLMo2222 by offering both theoretical and empirical perspectives within the Peri-LN scheme. Further discussion of the OLMo2222 is provided in Appendix G.
@@ 3.4 Stability Analysis in Normalization Strategies
We analyze training stability in terms of the magnitude of activation. To this end, we examine the gradient norm with respect to the weight of the final layer in the presence of massive activation. For the formal statements and detailed proofs, refer to Appendix D.
@@
(3) Post-LN (vanishing gradient). Consider the following sequence of operations:
$a=\mathrm{MLP}(x),o=x+a,\tilde{o}=\mathrm{Norm}(o),$ (7)
then
where $h:=\mathrm{ReLU}\left(xW^{(1)}+b^{(1)}\right)$. In this case, when a massive activation $\|h\|$ occurs, $\mathrm{Norm}(\cdot)$ introduces an overly suppressing factor $\|x+a\|$, which contains a separate huge residual signal $x$, potentially leading to a vanishing gradient $\|\partial\mathcal{L}/\partial W^{(2)}\|$.
We have compiled a Table 1 that provides a overview of the variance and gradient intuition for each layer normalization strategy.
Table 2: Average benchmark scores (with standard deviations) across $5$ different training seeds for Post-, Pre-, and Peri-Layer Normalization language models. Each model size excludes the embedding parameters. Loss denotes the evaluation loss on random samples of the C4444 dataset (Raffel et al., 2020). Arch. denotes architecture, and Avg. denotes the averaged benchmark score across tasks. SFT avg. denotes the averaged benchmark score across tasks of instruction fine-tuned models. When calculating the evaluation score, diverged checkpoints were excluded.

### v3

# Peri-LN: Revisiting Normalization Layer in the Transformer Architecture

@@ Abstract
Selecting a layer normalization (LN) strategy that stabilizes training and speeds convergence in Transformers remains difficult, even for today’s large language models (LLM). We present a comprehensive analytical foundation for understanding how different LN strategies influence training dynamics in large-scale Transformers. Until recently, Pre-LN and Post-LN have long dominated practices despite their limitations in large-scale training. However, several open-source models have recently begun silently adopting a third strategy without much explanation. This strategy places normalization layer peripherally around sublayers, a design we term Peri-LN. While Peri-LN has demonstrated promising performance, its precise mechanisms and benefits remain almost unexplored. Our in-depth analysis delineates the distinct behaviors of LN strategies, showing how each placement shapes activation variance and gradient propagation. To validate our theoretical insight, we conduct extensive experiments on Transformers up to $3.2$B parameters, showing that Peri-LN consistently achieves more balanced variance growth, steadier gradient flow, and convergence stability. Our results suggest that Peri-LN warrants broader consideration for large-scale Transformer architectures, providing renewed insights into the optimal placement of LN.
@@ 1 Introduction
Building on a rapidly expanding lineage of Transformer-based large language models, open-source models have shown remarkable impact (Hoffmann et al., 2022; Guo et al., 2025; Yoo et al., 2024). As the demand for larger and more powerful models grows, various training stabilization techniques have been introduced (Yang et al., 2022; Zhai et al., 2023; Loshchilov et al., 2024). Among these, the choice of where and how to apply layer normalization (LN: LayerNorm or RMSNorm; Ba et al., 2016; Zhang & Sennrich, 2019) critically influences model convergence (Xiong et al., 2020; Kedia et al., 2024; Wortsman et al., 2024). However, their immense computational requirements have restricted deeper exploration of the underlying Transformer structure. Are we truly employing the optimal LN placement? In practice, fully revealing the results of massive resource investments can be challenging (Rivière et al., 2024). Despite its importance, there is still no consensus on a single best LN placement strategy.
@@
Two prominent LN placements have been widely explored. Post-LN (Vaswani et al., 2017) normalizes the hidden state after adding the sub-layer output to the residual stream (that is, $\mathrm{Norm}(x+\mathrm{Module}(x))$ where $x$ is input hidden state. $\mathrm{Norm}$ is LN). This helps constrain the variance of hidden states but may inadvertently weaken gradient signals, particularly in deeper models (Kedia et al., 2024). Pre-LN (Dubey et al., 2024), by contrast, normalizes before passing the hidden state to the sub-layer (that is, $x+\mathrm{Module}(\mathrm{Norm}(x))$). While this can enhance gradient propagation, it also admits so-called “massive activations,” where hidden states grow exponentially across layers (Sun et al., 2024).
@@
Previous studies on deep convolutional neural networks (CNNs) have analyzed the impact of batch normalization on variance changes during the initialization stage of ResNet architectures, demonstrating its relationship to model performance (De & Smith, 2020). They noted that, in models without normalization, hidden activation growth at initialization can be exponential, leading to poor performance and stability. In contrast, in pre-normalized CNNs, the variance of hidden activations was shown to increase linearly as model depth grows. In the same vein, Kedia et al. (2024) reported that, for Transformer architectures as well, the variance in the forward propagation of Transformer-based language models at initialization increases linearly with depth. However, in the context of Transformer architectures, we observed that this variance growth at initialization does not persist as training progresses as shown in Figure 1. Sections 3, 4, 5, and 6 provide a more detailed discussion of these hidden-state growth patterns.
@@
Beyond these two common strategies, Post-LN and Pre-LN, a third LN placement has quietly emerged in large-scale open-source models: applying LN around the sub-layer, i.e., on both its input and output. Although recent open-source models (Team et al., 2025; Rivière et al., 2024; OLMo et al., 2024) have quietly adopted such designs and demonstrated promising performance on a large scale, these efforts often appeared isolated, lacking a conceptual unifying framework or a thorough investigation into their benefits. In this paper, we coin the term Peri-LN [^1] Peri- (Prefix) means “around,” reflecting that LN encapsulates the entire sub-layer. (e.g. peripherally) to unify these scattered approaches and highlight an underexplored avenue for stabilizing large-scale Transformer training. By dissecting the forward- and backward-pass dynamics of each LN strategy, we clarify how, when, and why they differ, interpreting these distinctions through the lens of training stability.
@@
    Present an in-depth analysis of Post-LN and Pre-LN in large-scale Transformers, examining how variance and gradient properties evolve beyond initialization.
@@
    Investigate Peri-LN to understand how normalizing both the inputs and outputs of each module moderates hidden-state behavior during forward and backward propagation, providing a systematic perspective on this underexplored alternative.
@@
    Provide quantitative evidence on how large activation influences training stability, benchmark performance, and model behaviors.
Figure 1: Illustration of hidden-state variance across different model depths and training iterations. From initialization through training on $6.3$ billion tokens, we observe the growth in hidden-state variance for both Pre-LN and Post-LN architectures. The analysis is based on a $1.5$B-parameter model. Detailed settings and additional results are provided in Section 5.1.
@@ 2 Background and Motivation
The analysis of activation variance at model initialization has long been central to understanding normalization layers and enhancing stability in convolutional neural networks (CNNs) (De & Smith, 2020; He et al., 2016; Brock et al., 2021a). Specifically, De & Smith (2020) showed that batch normalization in residual blocks can bias networks toward the identity function, thereby stabilizing gradients and improving overall training dynamics.
Similar investigations have emerged for Transformer architectures, examining how variance propagates and how gradients behave in both post-layer normalization (Post-LN) and pre-layer normalization (Pre-LN) configurations (Xiong et al., 2020; Kedia et al., 2024; Wortsman et al., 2024). Early work comparing Post- and Pre-LN primarily focused only on the initialization stage. Xiong et al. (2020) observed that Pre-LN architectures tend to exhibit more stable gradients, but can still encounter issues such as gradient spikes and divergence, especially in deeper models or large-scale pre-training scenarios (Zhai et al., 2023; Wortsman et al., 2024; Fishman et al., 2024; Chung et al., 2024).
Among these challenges, the phenomenon of “massive activations” has attracted attention (Dettmers et al., 2022; Yu et al., 2024; Fishman et al., 2024). Notably, Sun et al. (2024) identified that in Pre-LN architectures, large spikes in activation magnitude can persist across layers due to residual connections. These massive activations act as fixed biases, potentially narrowing the model’s focus to certain tokens and may influence generalization. However, the underlying mechanisms behind these large values, and their exact impact on the training process, remain not yet well understood.
@@
These observations underscore the ongoing need to clarify how activation dynamics and normalization strategies interact, especially in large-scale training.
In response, this work aims to deepen our understanding of how normalization strategies influence Transformer training, with particular attention to the emergence of large activations and their implications for stability and performance.
3 Normalization Strategies
In this section, we discuss how different placements of layer normalization (LN  [^2] Unless stated otherwise, LN refers to both LayerNorm (Ba et al., 2016) and RMSNorm (Zhang & Sennrich, 2019).) in Transformer architecture affect both training stability and the statistics of hidden states (activations  [^3] We use “hidden state” and “activation” interchangeably.).
@@ 3.2 Variance Behavior from Initialization to Training
During training, we find that Post-LN maintains a roughly constant variance, which helps avert exploding activations. However, as models grow deeper and training progresses, consistently normalizing $x_{l}+\mathrm{Module}(x_{l})$ can weaken gradient flow, occasionally causing partial vanishing gradients and slower convergence. In contrast, Pre-LN normalizes $x_{l}$ before the module but leaves the module output unnormalized, allowing hidden-state variance to accumulate exponentially once parameter updates amplify the input. Although Pre-LN preserves gradients more effectively in earlier stages, this exponential growth in variance can lead to “massive activations” (Sun et al., 2024), risking numeric overflow and destabilizing large-scale training.
Takeaways from Pre-LN & Post-LN.
(1) Keeping the Highway Clean: Post-LN’s Potential for Gradient Vanishing and Slow Convergence. When layer normalization is placed directly on the main path (Placement $C$ in Figure 2), it can cause gradient vanishing and introduce fluctuations in the gradient scale, potentially leading to instability (Xiong et al., 2020). (2) Maintaining a Stable Highway: Pre-LN May Not Suffice for Training Stability. Pre-LN does not normalize the main path of the hidden states, thereby avoiding the issues that Post-LN encounters. Nevertheless, a structural characteristic of Pre-LN is that any large values arising in the Attention or MLP modules persist through the residual identity path. In particular, as shown in Figure 1, the exponentially growing magnitude and variance of the hidden states in the forward path may lead to numerical instability.
3.3 Peri-Normalization in Transformers
Recent open-source Transformer architectures have placed normalization layers in unconventional placements (Rivière et al., 2024; Team et al., 2025; OLMo et al., 2024). In particular, these models apply an additional normalization layer at the module output (Output-LN), yet the benefits of this design choice remain unclear. To assess the impact of Output-LN, we analyze the Peri-LN architecture.
@@ 3.3 Peri-Normalization in Transformers
    (Optional) Initial Embedding Normalization:
@@
    $y_{l}=x_{l}+\mathrm{Norm}\Bigl{(}\mathrm{Module}\bigl{(}\mathrm{Norm}(x_{l})\bigr{)}\Bigr{)},$ (3)
@@
Both the latest Gemma (Team et al., 2025; Rivière et al., 2024) and OLMo (OLMo et al., 2024) model families, which apply output layer normalization, adopt the same peri-normalization strategy. However, neither line of work rigorously examines how this placement constrains variance or mitigates large residual activations. Our study extends these open-sourced large-scale models by providing both theoretical and empirical insights into the Peri-LN scheme.
@@
By normalizing both the input and output of each sub-layer, Peri-LN constrains the residual spikes commonly observed in Pre-LN, while maintaining a stronger gradient pathway than Post-LN. Concretely, if $\mathrm{Norm}(\mathrm{Module}(\mathrm{Norm}(x_{l})))$ exhibits near-constant variance $\beta_{0}$, then
$\mathrm{Var}(x_{l+1})\;\approx\;\mathrm{Var}(x_{l})+\beta_{0},$ (4)
resulting in linear or sub-exponential growth of activations, in contrast to the exponential growth patterns of Pre-LN.
Although Pre-LN and Peri-LN exhibit comparable, roughly linear variance growth at initialization (De & Smith, 2020; Xie et al., 2023), their trajectories diverge once training begins. The additional normalization layer (Output-LN) in Peri-LN preserves the conditions of Eq. 4, enabling the model’s hidden states to remain better conditioned. By contrast, the rapid surge in variance observed in Pre-LN can trigger instability during the early stages of training, an effect we quantify empirically in Sections 5 and 6.
@@
Figure 3: Performance comparison of Post-, Pre-, and Peri-LN Transformers during pre-training. Figure 3(a) llustrates the pre-training loss across learning rates. Pre-training loss and gradient norm of best performing $400$M size Transformers are in Figure 3(b) & 3(c).
(a) Divergence at seed $2$
(b) Loss spike at seed $3$
(c) Gradient spikes at seed $5$
(d) Loss spikes at seed $5$
Figure 4: Common case of early stage instability in pre-training. In most of our experiments across different random seeds, the Pre-LN architecture exhibited early-stage instability. Although we initially suspected that a high learning rate might be the root cause, lowering it did not substantially mitigate these issues. By contrast, under the same settings, Peri-LN displayed stable training curves.
@@ 3.4 Stability Analysis in Normalization Strategies
Xiong et al. (2020) showed that, at initialization, Pre-LN exhibits smaller gradient scales at the final layer compared to Post-LN, with respect to model depth. In this study, we broaden our analysis beyond initialization to monitor hidden state variance over the full course of training. Building on the earlier observation that the deepest layer exhibits the largest activations, we focus on this surge in the final layer under the Peri-LN strategy to clarify its impact on training stability. To this end, we analyze stability by examining the gradient norm with respect to the final layer weights in the presence of massive activation. Formal statements and detailed proofs are presented in Appendix C.
@@
The layer-wise amplification documented in §3.2, combined with the bounds in Proposition 3.1, naturally explains the gradient spikes, and occasional divergences that arise in Pre-LN during large-scale pre-training. We revisit this phenomenon in §4.3. By contrast, the additional normalization in Peri-LN acts as a self-regularizing mechanism that damps variance growth, making the architecture less sensitive to large activations and therefore more stable in practice. The formal analysis for Post-LN is deferred to Appendix B.
4 Experiments
In this section, we provide a comprehensive comparison of Post-, Pre-, and Peri-Layer Normalization (LN) across large-scale Transformer pre-training, instruction-tuning, and subsequent evaluations on the language domain.
Table 1: Average benchmark scores (with standard deviations) across $5$ different training seeds for Post-, Pre-, and Peri-Layer Normalization language models. Each model size excludes the embedding parameters. Loss denotes the evaluation loss on random samples of the C4444 dataset. Arch. denotes architecture, and Avg. denotes the averaged benchmark score across tasks. SFT avg. denotes the averaged benchmark score across tasks of instruction fine-tuned models. Diverged checkpoints are excluded from the evaluation score computation.